I need your help to solve for X & Y using the addition method.
4x-5y=7
-4x+5y=7

Answer by funmath(2873)

I need your help to solve for X & Y using the addition method.
Add the two equations and you get:
4x-5y=7
-4x+5y=7
_________
0+0=14
0=14
Both your variables are eliminated and 0 does not = 14. There is no solution.
This is an independent inconsistent system. Graphically, these would be two parallel lines that will never intersect to give you a solution.
Happy Calculating!!!